OpenStudy (anonymous):

trivia question 2x-5(x-3)=2(x-10)

OpenStudy (anonymous):

You forgot the L at the end of 'trivial.' :-p

OpenStudy (anonymous):

Simplify and then solve for x.

OpenStudy (anonymous):

I kid, I kid!

OpenStudy (anonymous):

i was about to say

OpenStudy (anonymous):

Use distributive property then start canceling stuff.

OpenStudy (anonymous):

i tried it but i couldnt figure it out

OpenStudy (anonymous):

the parenthesis get me confused

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Distributive property example: \(3(x+2)=3x+6\)

OpenStudy (anonymous):

\[2x-5(x-3)=2(x-10)\] \[2x^{2}-11x+15=2x-20\] \[2x^{2}-13x+35=0\]

OpenStudy (anonymous):

You should be able to finish it.

OpenStudy (anonymous):

How did you get an \(x^2\) out of that @Brittni0605 ?

OpenStudy (anonymous):

This is not a quadratic equation.

OpenStudy (anonymous):

and \(-11x\) ? I'm confused.

OpenStudy (anonymous):